Determining valuesConvertible bond  ​Craig's Cake Company has an outstanding issue of 20​-year convertible bonds with a $800 par value. These bonds are convertible into 90 shares of common stock. They have a 16​% annual coupon interest​ rate, whereas the interest rate on straight bonds of similar risk is 12​%.

a.  Calculate the straight bond value of this bond.

b. Calculate the conversion​ (or stock) value of the bond when the market price is ​$16 per share of common stock.

c.  What is the minimum market value of the​ bond?
Answer #1

Straight value of bond is equal to the present value of all future interest payments and the principal amount

= 800*16%*PVAF(12%, 20 years) + 800*PVF(12%, 20 years)

= 128*7.4694 +800*0.1037

= $1,039.0432

b)Conversion Value = Number of shares*Market price per share

= 90*16

= $1,440

c)Minimum Market value is the minimum of straight value and conversion value i.e. $1,039.0432